OpenStudy (anonymous):

for each of these expressions, apply the distributive property and the other number properties to write and simplify an equivalent expression. the question will be on the inside :)

OpenStudy (anonymous):

1. \[\frac{ 2a + 5 }{ 5 } 2. \[10\left(\begin{matrix}2a +5 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

Isn't number 1 already simplified? For number 2, you do 10/5 which is 2. then multiply 2 to (2a+5) and you get 4a+10

OpenStudy (anonymous):

\[2a\left(\begin{matrix}2a+5 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

Oh ok then multiply 2a to (2a+5) which equals 4a^2+10a then divide that all by 5 you can separate this into 2 fractions [(4a^2)/5]+(10a/5) you can simplify the second fraction by doing 10/5 [(4a^2)/5]+2a

OpenStudy (anonymous):

ok :) after that you'll get the answer??

OpenStudy (anonymous):

that is the simplified answer\[\frac{ 4a^{2} }{ 5 }+2a\]

OpenStudy (anonymous):

ok :D \[2\left(\begin{matrix}2a^2+5+8 \\ 5a\end{matrix}\right)\]

OpenStudy (anonymous):

simplify the numerator and you get (2a^2+13). multiply 2 to the numerator and you get \[4a^{2}+26\] now divide that by 5a\[\frac{ 4a^{2}+26 }{ 5 }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

are you finding the area?

OpenStudy (anonymous):

yes